how many ounces of a 22% alcohol solution and a 43% alcohol solution must be combined to obtain 33 ounces of a 29% solution

Answer by ewatrrr(24785)

 
Hi,
Let x and (33-x) be the amounts of 43% and 22%
.43x + .22(33-x) = .29*33
43x + 22(33-x) = 29*33
11x = 7*33
x = 21 ounces of 43% and 11 ounces of 22%